Capital Rationing Decision for a Service Company Involving Four Proposals

Clearcast Communications Inc. is considering allocating a limited amount of capital investment funds among four proposals. The amount of proposed investment, estimated Operating income, and net cash flow for each proposal are as follows:

Investment Year Operating Income Net Cash Flow
Proposal A: $450,000 1 $30,000 $120,000
2    30,000   120,000
3    20,000   110,000
4    10,000   100,000
5   (30,000)    60,000
$60,000 $510,000
Proposal B: $200,000 1 $60,000 $100,000
2    40,000     80,000
3    20,000     60,000
4   (10,000)     30,000
5 (20,000)     20,000
$90,000 $290,000
Proposal C: $320,000 1 $36,000 $100,000
2    26,000    90,000
3    26,000    90,000
4    16,000    80,000
5    16,000    80,000
$120,000 $440,000
Proposal D: $540,000 1 $92,000 $200,000
2   72,000   180,000
3    52,000   160,000
4    12,000   120,000
5     (8,000)   100,000
$220,000 $760,000

The company's capital rationing policy requires a maximum cash payback period of three years. In addition, a minimum average rate of return of 12% is required on all projects. If the preceding standards are met, the net present value method and present value indexes are used to rank the remaining proposals.

Present Value of $1 at Compound Interest
Year 6% 10% 12% 15% 20%
1 0.943 0.909 0.893 0.870 0.833
2 0.890 0.826 0.797 0.756 0.694
3 0.840 0.751 0.712 0.658 0.579
4 0.792 0.683 0.636 0.572 0.482
5 0.747 0.621 0.567 0.497 0.402
6 0.705 0.564 0.507 0.432 0.335
7 0.665 0.513 0.452 0.376 0.279
8 0.627 0.467 0.404 0.327 0.233
9 0.592 0.424 0.361 0.284 0.194
10 0.558 0.386 0.322 0.247 0.162

Required:

1. Compute the cash payback period for each of the four proposals.

Cash Payback Period
Proposal A: 4 years
Proposal B: 2 years, 4 months
Proposal C: 3 years, 6 months
Proposal D: 3 years

2. Giving effect to straight-line depreciation on the investments and assuming no estimated residual value, compute the average rate of return for each of the four proposals. If required, round your answers to one decimal place.

Average Rate of Return
Proposal A: %
Proposal B: %
Proposal C: %
Proposal D: %

3. Using the following format, summarize the results of your computations in parts (1) and (2). By placing the calculated amounts in the first two columns on the left and indicate which proposals should be accepted for further analysis and which should be rejected. If required, round your answers to one decimal place.

Proposal Cash Payback Period Average Rate of Return Accept or Reject
A 4 years % Reject
B 2 years, 4 months % Accept
C 3 years, 6 months % Reject
D 3 years % Accept

4. For the proposals accepted for further analysis in part (3), compute the net present value. Use a rate of 12% and the present value of $1 table above. Round to the nearest dollar.

Select the proposal accepted for further analysis. Proposal B Proposal D
Present value of net cash flow total $ $
Amount to be invested $ $
Net present value $ $

5. Compute the present value index for each of the proposals in part (4). If required, round your answers to two decimal places.

Select proposal to compute Present value index. Proposal B Proposal D
Present value index (rounded)
Answer #1

1. cash payback period

Investment Year cummulative Cash Flow
Proposal A: $450,000 1 $120,000
2 240000 [120000+120000]
3 350000 [240000+110000]
4 450000
5    510000

initial investment will covered in 4th year

so payback period is 4years

cummulative cash flow
Proposal B: $200,000 1 $100,000
2     1,80,000
3     240000 [180000+60000]
4     270000
5     290000

year before cash recovery + unrecovered cash/ cash flow in next period

=2 + [200000-180000]/60000

=2+0.33 =2+ (12months*0.33)

=2 year 4months

cummulative cash flow
Proposal C: $320,000 1 $100,000
2 1, 90,000
3    280000[190000+90000]
4    360000 [280000+80000]
5    440000 [360000+80000]

= 3 + [320000-280000]/80000

=3 years 6 months

Proposal D: $540,000 1 $200,000
2 380,000[200000+180000]
3 540000
4 660000
5 760000

3 years 540000 initial outflow will be covered back.

2]average rate of return = average income/ average investment

A = 60000/5=12000 average income

investment = 450000+0salvage/2

=225000

average rate of return = 12000/225000

=5.33%

B =90000/5 ] / [200000/2]

=18000/100000

=18%

C=[120000/5]/[320000/2]

=24000/160000

=15%

D=[220000/5]/[540000/2]

=44000/270000

=16.30%

3]

Proposal Cash Payback Period Average Rate of Return Accept or Reject
A 4 years 5.33 % Reject
B 2 years, 4 months 18 % Accept
C 3 years, 6 months 15 % Reject
D 3 years 16.30 % Accept

4]

Select the proposal accepted for further analysis. Proposal B Proposal D
Present value of net cash flow total $226200 $569000
Amount to be invested $(200000) $(540000)
Net present value $26200 $29000
PV FACTOR NET CASH FLOW PRESENT VALUE OF CASH FLOW
Proposal B: $200,000 1 0.893 $100,000 89300
2 0.797     80,000 63760
3 0.712     60,000 42720
4 0.636     30,000 19080
5 0.567     20,000 11340
PRESENT VALUE 226200
PV FACTOR NET CASH FLOW PRESENT VALUE
Proposal D: $540,000 1 0.893 $200,000 178600
2 0.797   180,000 143460
3 0.712   160,000 113920
4 0.636   120,000 76320
5 0.567   100,000 56700
TOTAL 569000

5.PROFITABILITY INDEX = PRESENT VALUE OF NET CASH INFLOW/ CASH OUTFLOW

B =226200/200000

=1.13

D= 569000/540000

=1.05